# Coloring Hardness on Low Twin-Width Graphs

### v1

@@ 1 Introduction
We also show that 3-Coloring is $\mathsf{NP}$-hard on $\mathcal{T}_{4}$. Note that planar graphs have twin-width at least 7 [33] (and at most 8 [27]). So we cannot just invoke the hardness of 3-Coloring in planar graphs. One would need to provide 4-sequences for the graphs produced by the reduction. Another option would be to tune long subdivisions and turn them into hard 3-Coloring instances. Instead, we give an ad hoc construction for which both the correctness and membership in $\mathcal{T}_{4}$ are easy to verify.
@@
Another consequence of the abovementioned result of [2] is that $\mathcal{T}_{3}$ excludes a fixed planar graph as an induced minor. Thus Questions˜1.7 and 1.6 are special cases of the same questions on any class excluding a planar induced minor, which was previously raised for Max Independent Set [18]. The class $\mathcal{T}_{3}$ is a good candidate for a negative answer to the latter question (on the pessimistic side), or Question˜1.7 could serve as a preliminary step in positively answering it (on the optimistic side). There are other questions and conjectures on classes excluding a planar induced minor. They can be revisited on the particular class $\mathcal{T}_{3}$, like for instance the following conjecture appearing in [26].
@@ 3 Hardness of Coloring Graphs of Twin-Width 3
As a decision problem, Min Coloring inputs a graph $G$ and an integer $k$ and asks whether the chromatic number of $G$, $\chi(G)$, is at most $k$. As a function problem, one is given the mere graph $G$ and has to output a (proper) $\chi(G)$-coloring of $G$. We show that Min Coloring is $\mathsf{NP}$-hard on $\mathcal{T}_{3}$, even in its decision form and when a 3-sequence of $G$ is given in input. As Min Coloring is polynomial-time solvable on cographs, on trees, and on unit circular-arc graphs (all of which have twin-width at most 3), and planar graphs are not included in $\mathcal{T}_{3}$, this requires finding a novel kind of encoding of rich structures (here, 3-Sat instances) onto graphs of twin-width at most 3.
@@ 4 Hardness of 3-Coloring Graphs of Twin-Width 4
3-Coloring is $\mathsf{NP}$-hard and, unless the ETH fails, requires $2^{\Omega(\sqrt{N})}$ time on $N$-vertex graphs of twin-width at most 4 (even if a 4-sequence is provided).
@@
We reduce from the $\mathsf{NP}$-complete Not-All-Equal 3-Sat problem [35] (where each clause is on three distinct variables). Furthermore, by the Sparsification Lemma [29], unless the ETH fails, there is no $2^{o(n)}$-time algorithm for $n$-variable $O(n)$-clause instances. Indeed, the reduction from 3-Sat to Not-All-Equal 3-Sat in [35] is linear. Let $x_{1},\ldots,x_{n}$ be the variables of an instance $\varphi$ and $c_{1},\ldots,c_{m}$ its clauses, with $m=O(n)$. We build a graph $G:=G(\varphi)$ as follows.
For each clause $c_{j}=\ell_{1}\lor\ell_{2}\lor\ell_{3}$, add a triangle $t_{j}$ on three vertices $u_{j},v_{j},w_{j}$, where $u_{j},v_{j},w_{j}$ correspond to $\ell_{1},\ell_{2},\ell_{3}$, respectively.
For each variable $x_{i}$, start with the $m$-vertex path $x_{i,1}x_{i,2}\cdots x_{i,m}$. Let $j_{1}<j_{2}<\cdots<j_{q}$ be the clause indices in which $x_{i}$ appears, with $q\geqslant 1$. For each $h\in[q-1]$, let $s_{h},s_{h+1}\in\{\neg,\varepsilon\}$ be the signs of $x_{i}$ in clauses $c_{j_{h}}$ and $c_{j_{h+1}}$. If either
  • $|j_{h+1}-j_{h}|$ is even and $s_{h}\neq s_{h+1}$, or
  • $|j_{h+1}-j_{h}|$ is odd and $s_{h}=s_{h+1}$,
subdivide once the edge $x_{i,j_{h+1}-1}x_{i,j_{h+1}}$ and denote the new vertex by $x^{\prime}_{i,j_{h+1}}$. Let $P_{i}$ be the resulting (still induced) path, which we refer to as the variable path (of $x_{i}$).
We add one vertex $z$ adjacent to every vertex of $\bigcup_{i\in[n]}V(P_{i})$. Finally, for each clause $c_{j}=s_{1}x_{i_{1}}\lor s_{2}x_{i_{2}}\lor s_{3}x_{i_{3}}$, we add the edges $u_{j}x_{i_{1},j}$, $v_{j}x_{i_{2},j}$, and $w_{j}x_{i_{3},j}$ (regardless of $s_{1},s_{2},s_{3}$). This concludes the construction; see Figure˜3 for an illustration.
[…]
Figure 3: The graph $G(\varphi)$ for the 7-variable 8-clause formula $\varphi$ with: $c_{1}=x_{1}\lor\neg x_{2}\lor x_{3}$, $c_{2}=\neg x_{1}\lor x_{4}\lor x_{5}$, $c_{3}=x_{2}\lor\neg x_{3}\lor x_{6}$, $c_{4}=x_{1}\lor x_{6}\lor\neg x_{7}$, $c_{5}=x_{4}\lor x_{5}\lor x_{7}$, $c_{6}=x_{2}\lor x_{4}\lor\neg x_{6}$, $c_{7}=\neg x_{1}\lor\neg x_{5}\lor x_{7}$, $c_{8}=x_{3}\lor\neg x_{6}\lor\neg x_{7}$.
Note that $G$ has $N$ vertices with $N\leqslant 3m+(2m-1)n+1=O(n^{2})$. As we will next show that the reduction is correct and the twin-width of $G$ is at most 4, this establishes the claimed ETH lower bound. Indeed, deciding if $G$ is 3-colorable in $2^{o(\sqrt{N})}$ time would solve the Not-All-Equal 3-Sat instance $\varphi$ in $2^{o(n)}$ time.
Parity property along the variable paths. Fix a variable $x_{i}$ with occurrences at indices $j_{1}<\cdots<j_{q}$ and signs $s_{1},\ldots,s_{q}$, respectively. By construction, for every $h\in[q-1]$, the distance in $P_{i}$ between $x_{i,j_{h}}$ and $x_{i,j_{h+1}}$ is even if and only if $s_{h}=s_{h+1}$. Consequently, in any (proper) $2$-coloring of $P_{i}$,
$c(x_{i,j_{h}})=c(x_{i,j_{h^{\prime}}})\Longleftrightarrow s_{h}=s_{h^{\prime}},\quad\text{for all }h,h^{\prime}\in[q],$
i.e., all positive occurrences receive the same color and all negative occurrences receive the other color. Indeed, an even distance between $x_{i,j_{h}}$ and $x_{i,j_{h+1}}$ (thus with equal color in a $2$-coloring of $P_{i}$) does not change the sign of $x_{i}$, whereas an odd distance flips the sign of $x_{i}$.
If $\bm{\varphi}$ is satisfiable, then $\bm{G}$ is $3$-colorable. Let $\mathcal{A}$ be a truth assignment such that every clause of $\varphi$ has at least one literal set to true and at least one literal set to false. We define a 3-coloring $c$ of $G$. First, set $c(z)=3$.
For each variable $x_{i}$, pick (arbitrarily) one clause index $j^{*}$ such that $x_{i}$ appears in $c_{j^{*}}$, and let $s^{*}$ be the sign of $x_{i}$ in $c_{j^{*}}$. Set $c(x_{i,j^{*}})=1$ if $\mathcal{A}$ sets literal $s^{*}x_{i}$ to true, and $c(x_{i,j^{*}})=2$ otherwise. Since $z$ has color $3$, the path $P_{i}$ must use only colors in $\{1,2\}$, and the coloring of $P_{i}$ is uniquely determined by the color of $x_{i,j^{*}}$. By the parity property above, for every occurrence of $x_{i}$ in a clause $c_{j}$ with sign $s$, the vertex $x_{i,j}$ has color $1$ if and only if $\mathcal{A}$ sets $sx_{i}$ to true (and color $2$ otherwise).
Now fix a clause $c_{j}=\ell_{1}\lor\ell_{2}\lor\ell_{3}$ with triangle $t_{j}=\{u_{j},v_{j},w_{j}\}$. The three neighbors $x_{i_{1},j},x_{i_{2},j},x_{i_{3},j}$ of $t_{j}$ are colored in $\{1,2\}$ and, since $\mathcal{A}$ satisfies $\varphi$, these three colors are not all equal. Then, let $a,b\in[3]$ be such that $c(x_{i_{a},j})\neq c(x_{i_{b},j})$. Give the neighbor of $x_{i_{a},j}$ (resp. $x_{i_{b},j}$) in $t_{j}$ the color of $\{1,2\}$ opposite to $c(x_{i_{a},j})$ (resp. $c(x_{i_{b},j})$). Give the third vertex of $t_{j}$ color 3. This yields a proper $3$-coloring of $t_{j}$ extending the already fixed colors on the variable paths. Doing this independently for all clauses completes a (proper) $3$-coloring of $G$; see Figure˜4.
[…]
Figure 4: A 3-coloring of $G(\varphi)$ for the formula $\varphi$ of Figure˜3 corresponding to the assignment where all variables are set to true except $x_{4}$, which is set to false.
If $\bm{G}$ is $3$-colorable, then $\bm{\varphi}$ is satisfiable. Let $c$ be a (proper) $3$-coloring of $G$. Up to color permutation, we can assume that $c(z)=3$. Since $z$ is adjacent to every vertex of every $P_{i}$, each path $P_{i}$ uses only colors in $\{1,2\}$.
We define an assignment $\mathcal{A}$ of the variables of $\varphi$ as follows. For each variable $x_{i}$, if $x_{i}$ appears positively in some clause $c_{j}$, $\mathcal{A}$ sets $x_{i}$ to true if $c(x_{i,j})=1$, and to false otherwise; if $x_{i}$ appears only negatively, $\mathcal{A}$ sets $x_{i}$ to false if $c(x_{i,j})=1$ for some clause $c_{j}$ containing literal $\neg x_{i}$, and to true otherwise. This is well-defined by the parity property: all positive occurrences have the same color and all negative occurrences have the other color (in $\{1,2\}$).
Consider any clause $c_{j}=\ell_{1}\lor\ell_{2}\lor\ell_{3}$ represented by triangle $t_{j}$. If the three literals $\ell_{1},\ell_{2},\ell_{3}$ had the same truth value under $\mathcal{A}$, then the three neighbors of $t_{j}$ on the variable paths would all receive the same color in $\{1,2\}$. In that case, each of $u_{j},v_{j},w_{j}$ would be forbidden to use that color, leaving only two colors for the triangle $t_{j}$, which would not result in a 3-coloring of $G$. Hence, in every clause, not all three literals have the same truth value under $\mathcal{A}$. Therefore, $\mathcal{A}$ satisfies $\varphi$ (as a Not-All-Equal 3-Sat instance).
$\bm{G}$ admits a $4$-sequence. We describe a partition sequence for $G$ of width at most 4. It is computable in polynomial time and does not depend on $\varphi$ beyond knowing $n$ and $m$. The sequence has four stages.
In the first stage, we merge each clause gadget into a single part. For each $j\in[m]$, merge $\{u_{j}\}$ with $\{v_{j}\}$, then merge the result with $\{w_{j}\}$. We denote the resulting part by $T_{j}:=\{u_{j},v_{j},w_{j}\}$. Throughout this stage, every non-singleton part has red degree at most 3, and every singleton part has at most one red neighbor.
In the second stage, we merge every existing $x^{\prime}_{i,j}$ with $x_{i,j}$, and set $X_{i,j}:=\{x^{\prime}_{i,j},x_{i,j}\}$ if $x^{\prime}_{i,j}$ exists, and $X_{i,j}:=\{x_{i,j}\}$ otherwise. This does not create any part of red degree at least 4. The rest of the partition sequence has width at most 4, even if every variable path (which has now exactly $m$ vertices) is fully red. We thus assume that this is the case. Henceforth, for each $j\in[m]$, we keep denoting by $X_{1,j}$ the part containing $x_{1,j}$.
In the third stage, for $i$ going from 2 to $n$, for $j$ going from 1 to $m$, we merge $X_{1,j}$ with $X_{i,j}$. This merge cannot increase the red degree of a part that is not the resulting part (recall that we assumed that the variable paths are fully red), and the latter has at most four red neighbors: $T_{j}$ (if it is adjacent to), $X_{1,j-1}$ (if it exists), $X_{1,j+1}$ and $X_{i,j+1}$ (if they exist). Indeed note that at this point, $X_{i,j-1}$ has already been merged with $X_{1,j-1}$.
In the last stage, for $j$ going from 2 to $m$, we merge $X_{1,1}$ and $X_{1,j}$, and keep denoting the resulting part by $X_{1,1}$, and we merge $T_{1}$ and $T_{j}$, and keep denoting the resulting part by $T_{1}$. This never creates a part with red degree larger than 3. Finally, there are three parts left: $T_{1},X_{1,1},\{z\}$, which we merge in any order. The concatenation of the four stages is a 4-sequence for $G$, which concludes the proof.
@@
For every $\geqslant 3, kk$-Coloring is $\mathsf{NP}$-hard and, unless the ETH fails, requires $2^{\Omega(\sqrt{n})}$ time on $n$-vertex graphs of twin-width at most 4.
@@
Adding k3k-3 universal vertices to the graph $G(\varphi)$ built in the proof of Theorem˜1.2 does not change its twin-width and makes the resulting graph $k$-colorable if and only if $G(\varphi)$ is 3-colorable.

### v2

@@ 1 Introduction
We also show that 3-Coloring is $\mathsf{NP}$-hard on $\mathcal{T}_{4}$. Note that planar graphs have twin-width at least 7 [34] (and at most 8 [28]). So we cannot just invoke the hardness of 3-Coloring in planar graphs. We tune long subdivisions to turn them into hard 3-Coloring instances.
@@
Another consequence of the above result of [2] is that $\mathcal{T}_{3}$ excludes a fixed planar graph as an induced minor. Thus Questions˜1.7 and 1.6 are special cases of the same questions on any class excluding a planar induced minor, which was previously raised for Max Independent Set [19]. The class $\mathcal{T}_{3}$ is a good candidate for a negative answer to the latter question (on the pessimistic side), or Question˜1.7 could serve as a preliminary step in positively answering it (on the optimistic side). There are other questions and conjectures on classes excluding a planar induced minor. They can be revisited on the particular class $\mathcal{T}_{3}$, like for instance the following conjecture appearing in [27].
@@ 3 Hardness of Coloring Graphs of Twin-Width 3
As a decision problem, Min Coloring takes as input a graph $G$ and an integer $k$ and asks whether the chromatic number of $G$, $\chi(G)$, is at most $k$. As a function problem, one is given the mere graph $G$ and has to output a (proper) $\chi(G)$-coloring of $G$. We show that Min Coloring is $\mathsf{NP}$-hard on $\mathcal{T}_{3}$, even in its decision form and when a 3-sequence of $G$ is given as input. As Min Coloring is polynomial-time solvable on cographs, on trees, and on unit circular-arc graphs (all of which have twin-width at most 3), and planar graphs are not included in $\mathcal{T}_{3}$, this requires finding a novel kind of encoding of rich structures (here, 3-Sat instances) onto graphs of twin-width at most 3.
@@ 4 Hardness of 3-Coloring Graphs of Twin-Width 4
3-Coloring is $\mathsf{NP}$-hard and, unless the ETH fails, requires $2^{\Omega(n/\log n)}$ time on $n$-vertex graphs of twin-width at most 4 (even if a 4-sequence is provided).
@@
It is known that 3-Coloring is $\mathsf{NP}$-hard on $n$-vertex graphs of maximum degree at most 4, and requires $2^{\Omega(n)}$ time, unless the ETH fails (see for instance [18, Lemma 1]).
Let $G$ be an $n$-vertex graph of degree at most 4. We build an $O(n\log n)$-vertex graph $G^{\prime}$ of twin-width at most 4 such that $G^{\prime}$ is 3-colorable if and only if $G$ is 3-colorable. Let $q:=2\lceil\log n\rceil$. Graph $G^{\prime}$ is obtained from $G$ by subdividing each edge $e\in E(G)$ $q$ times, thereby creating vertices $v_{e,1},\ldots,v_{e,q}$, and then adding one true twin (i.e., an adjacent vertex with the same closed neighborhood) $v^{\prime}_{e,1},v^{\prime}_{e,3},v^{\prime}_{e,5},\ldots,v^{\prime}_{e,q-1}$ to $v_{e,1},v_{e,3},v_{e,5},\ldots,v_{e,q-1}$, respectively (that is, every other created vertex gets a true twin). The “orientation” of $e$ to fully specify this process is irrelevant, so we leave it undefined. (To get a unique construction, one can totally order $V(G)$, and always “orient” the edge from its smaller endpoint to its larger endpoint.)
The way $G^{\prime}$ is built, it can be observed that there is a partial 0-sequence from $G^{\prime}$ to the $q$-subdivision of $G$. Since every $(\geqslant 2\log n)$-subdivision of an $n$-vertex graph has twin-width at most 4 [5], $G^{\prime}$ has twin-width at most 4. The number of vertices of $G^{\prime}$ is $O(n+|E(G)|\log n)=O(n\log n)$, and its construction can be done in $O(n\log n)$ time.
Given an original vertex $u$ (i.e., $u\in V(G)$) of $G^{\prime}$ adjacent to some $v_{e,1}$, and a color $i\in[3]$, the 3-coloring propagation of $u$ along $e$ is the only (proper) 3-coloring of
$G[\{u,v_{e,1},v^{\prime}_{e,1},v_{e,2},v_{e,3},v^{\prime}_{e,3},\ldots,v_{e,q-1},v^{\prime}_{e,q-1},v_{e,q}\}]$
that gives color $i$ to $u$. Note that this 3-coloring also gives color $i$ to $v_{e,q}$.
If $G$ is 3-colorable, we start by copying a 3-coloring of $G$ at the original vertices of $G^{\prime}$. We then make the 3-coloring propagation for each pair $u\in V(G),v_{e,1}$ with $uv_{e,1}\in E(G^{\prime})$. By the previous remark, this defines a (proper) 3-coloring of $G^{\prime}$. Conversely, the projection of any 3-coloring of $G^{\prime}$ onto $V(G)$ properly colors $G$.
@@
For every $\geqslant 3, kk$-Coloring is $\mathsf{NP}$-hard and, unless the ETH fails, requires $2^{\Omega(n/\log n)}$ time on $n$-vertex graphs of twin-width at most 4.
@@
Adding k3k-3 universal vertices to the graph $G^{\prime}$ built in the proof of Theorem˜1.2 does not change its twin-width and makes the resulting graph $k$-colorable if and only if $G^{\prime}$ is 3-colorable.